Based on the following information, what is the standard deviation of returns?

State of Economy Probability of State
of Economy
Rate of Return if
State Occurs
Recession .23 .091
Normal .46 .106
Boom .31 .216
Answer #1

First we will calculate the mean return as per below:

The formula for mean is:

Mean = p1 * r1 + p2 * r2 + p3 * r3

where, p1,p2 and p3 are the probabilities and r1,r2 and r3 are the returns.

Putting the given values of the probability in the above formula, we get,

Mean = (0.23 * -.091%) + (0.46 * .106%) + (0.31 * .216%)

Mean = -.02093 + 0.04876 + 0.06696

Mean = 0.09479

Now, steps for calculating standard deviation are:

First we will calculate the deviation of returns from the mean return as per below:

Recession : -0.091 - 0.09479 = 0.18579

Normal : 0.106 - 0.09479 = 0.01121

Boom 0.216 - 0.09479 = 0.12121

In the next step, we will square the deviations computed above, as per below:

Recession: (0.18579)2 = 0.0345179241

Normal: (0.01121)2 = 0.0001256641

Boom : (0.12121)2 = 0.0146918641

In the next step, we will multiply the squared deviations computed above with their probabilities as per below:

Recession: 0.0345179241 * 0.23 = 0.007939122543

Normal:  0.0001256641 * 0.46 = 0.000057805486

Boom: 0.0146918641 * 0.31 = 0.004554477871

In the next step we will add up the values calculated above to find the variance, as per below:

Variance = 0.007939122543 + 0.000057805486 + 0.004554477871 = 0.01255141

In the final step, we will square root the variance calculated above to find the standard deviation:

Standard deviation = (0.01255141)1/2 = 0.11203306